A container is in the shape of a right cicular cone. The base area is 4.8 square feet and the height is 1.5 feet. Compute the capacity of the container in gallons.
Answer by checkley77(12844)

V=BH/3
V=4.8*1.5=7.2 FEET^3.
[1 GALLON=3.7853 LITERS & 1 LITER=61.0234 IN^3]
1 GALLON=3.7853*61.0234=230.99 IN^3
7.2FT^3=7.2*12^3=7.2*1728=12,441.6 IN^3
12,441.6/230.99=53.86 GALLONS.